A 3.80 kg fish swims upstream against the current. In order to make it over a rock jutting out of the water, the fish needs to accelerate upwards at an angle and then jump the rock barrier. The fish needs to both overcome the horizontal force of the current, which is 131 N, as well as make it over the rocks with a vertical force of 97.0 N. What is the acceleration of the fish just before it makes its jump? 


1
Expert's answer
2021-02-10T15:20:43-0500
"a=\frac{F}{m}\\a=\frac{\sqrt{F_x^2+F_y^2}}{m}\\a=\frac{\sqrt{131^2+97^2}}{3.8}=42.9\frac{m}{s^2}"